OpenStudy (anonymous):

A four- digit number ( with no repetitions) is to be formed from the set of number ( 0,1,2,3,4,5,6,7). Find the probability that the number : a) is even

Parth (parthkohli):

You first find the number of four-digit numbers that may be formed from the above set.

OpenStudy (anonymous):

Think of it this way, pick the numbers from right to left, least significant first.

Parth (parthkohli):

Can you do that? \[8P4 \] But there is something—we can't place a zero.

OpenStudy (mihai):

the probability is 50 %

OpenStudy (anonymous):

i did 6x6x5x4 ------- 7x7x6x5 ?

Parth (parthkohli):

Let's think of slots. _ _ _ _ Four slots. We have 7 choices for the first(there are eight numbers in the set, but we can't place a zero). 7 for the second too(we have a zero with us the second time). 6 for the third. (We already have two numbers filled out of eight). 5 for the fourth. (We have three filled).

Parth (parthkohli):

We multiply \(7 \times 7 \times 6 \times 5 \Longrightarrow 49 \times 30 \)

OpenStudy (anonymous):

yes then? number of even numbers = 6x6x5x4 i that correct

Parth (parthkohli):

First find the number of four digit numbers you could make, then divide that by 2.

OpenStudy (anonymous):

um answer is not 1/2 though

OpenStudy (mihai):

every number that ends with a 0,2,4 and 6 is an even number, so if 4 out of 8 digits give you an even number wouldn't that be 50/50?

OpenStudy (anonymous):

@ParthKohli see my post above. Make sense?

Parth (parthkohli):

Yup! That's correct!

OpenStudy (anonymous):

@ParthKohli answer is given as 780/1470 not 735/1470?